OpenStudy (rayman5766):

A $5000 investment earns $1575 simple interest in three years. What is the annual interest rate? A. 1.05% B. 3.15% C. 10.5% D. 31.5%

OpenStudy (igreen):

Use the formula: I = prt 1575 = (5000)(3)r Multiply 5000 * 3

OpenStudy (rayman5766):

is my answer 10.5

OpenStudy (igreen):

Correct.

OpenStudy (igreen):

500 * 3 = 15000 1575 = 15000r 1575 / 15000 = 0.105 0.105 * 100 = 10.5% r = 10.5%

OpenStudy (anonymous):

how did you get the 3?